OpenStudy (anonymous):

Find the difference. 8x - (12x2 + 5x) = Select one of the options below as your answer: A. -9x B. -12x2 + 3x C. -12x2 + 5x D. 12x2 + 3x

OpenStudy (munish):

8X-(12X2+5X) =8X-12X2-5X =-12X2+8X-5X =-12X2+3X so B)